# INTER: Mitigating Hallucination in Large Vision-Language Models by Interaction Guidance Sampling

### v1

@@ Abstract
Hallucinations in large vision-language models (LVLMs) pose significant challenges for real-world applications, as LVLMs may generate responses that appear plausible yet remain inconsistent with the associated visual content. This issue rarely occurs in human cognition. We argue that this discrepancy arises from humans’ ability to effectively leverage multimodal interaction information in data samples. Specifically, humans typically first gather multimodal information, analyze the interactions across modalities for understanding, and then express their understanding through language. Motivated by this observation, we conduct extensive experiments on popular LVLMs and obtained insights that surprisingly reveal human-like, though less pronounced, cognitive behavior of LVLMs on multimodal samples. Building on these findings, we further propose INTER: Interaction Guidance Sampling, a novel training-free algorithm that mitigate hallucinations without requiring additional data. Specifically, INTER explicitly guides LVLMs to effectively reapply their understanding of multimodal interaction information when generating responses, thereby reducing potential hallucinations. On six benchmarks including VQA and image captioning tasks, INTER achieves an average improvement of up to 3.4% on five LVLMs compared to the state-of-the-art decoding strategy. The code will be released when the paper is accepted.
@@ 3.5 Verifying the Impact of Interactions
It is challenging to quantify the relationship between the strength of multimodal interactions for keyword tokens and the overall accuracy of the response. To address this, we propose conducting experiments on the MME [19] benchmark for convenient verification. In MME, the keywords in the responses are limited to “yes” or “no”, which are also used to determine the accuracy of the responses. This setup allows us to measure the strength of multimodal interactions for keyword tokens using $\mathrm{I}(A)_{r}^{\textbf{y}_{t}}$, where $\textbf{y}_{t}\in\{\textit{yes},\textit{no}\}$ and $A=\{\textbf{v},\textbf{p}\}$, for each response $r$. The appearance of the yes/no keywords directly indicates the binary correctness of the response. Furthermore, to better verify the trend of the positive impact of multimodal interactions, we propose to partition all the generated responses on MME into multiple bins based on the value of $\mathrm{I}(A)_{r}^{\textbf{y}_{t}}$ ($\textbf{y}_{t}\in\{yes,no\}$) sorted in ascending order, with each bin containing $Q$ samples. Then, we calculate the average contribution of interactions within each bin $\mathbf{b}$, denoted by $\overline{I}_{\mathbf{b}}=\frac{1}{Q}\sum_{r\in\mathbf{b}}\mathrm{I}(A)_{r}^{\textbf{y}_{t}}$. We denote $\overline{C}_{\mathbf{b}}$ as the ratio of correct responses in each bin $\mathbf{b}$. Thus, we can analyze the correlation between $\overline{I}_{\mathbf{b}}$ and $\overline{C}_{\mathbf{b}}$ among different bins, allowing us to verify trends in the positive impact of multimodal interactions on response accuracy.
@@ 5.2 Experimental Results
Datasets Order InstructBLIP [13] LLaVA-v1.5 [38] Qwen-VL [3] mPLUG-owl2 [62]
Mean Absolute Value Range Mean Absolute Value Range Mean Absolute Value Range Mean Absolute Value Range
MME [19] $\mathrm{I}\left(A\right|\{\textbf{v},\textbf{p}\})^{\textbf{y}_{t}}$ 0.80 [-9.8, 5.7] 0.59 [-13.8, 5.3] 2.10 [-16.1, 10.0] 0.07 [-20.4, 27.8]
$\mathrm{I}\left(A\right|\{\textbf{p},\textbf{v}\})^{\textbf{y}_{t}}$ 4.07 [-7.4, 22.8] 0.59 [-13.8, 9.3] 1.93 [-10.6, 8.2] 0.07 [-16.5, 10.2]
CHAIR [46] $\mathrm{I}\left(A\right|\{\textbf{v},\textbf{p}\})^{\textbf{y}_{t}}$ 0.60 [-16.5, 14.2] 0.06 [-3.3, 3.9] 0.56 [-17.8, 12.4] 0.16 [-8.0, 7.7]
$\mathrm{I}\left(A\right|\{\textbf{p},\textbf{v}\})^{\textbf{y}_{t}}$ 0.60 [-3.3, 3.9] 0.06 [-3.3, 3.9] 0.94 [-25.4, 8.1] 0.16 [-7.9, 7.5]
Table 2: The range of the metric $\mathrm{I}(A)^{\textbf{y}_{t}}$.
[…]
$C_{s}$$\downarrow$ $C_{I}$$\downarrow$
InstructBLIP [13] M3ID [17] 63.1 21.1 29.8 1440.6
Ritual [59] 62.1 20.9 29.5 1576.7
SID [27] 59.7 21.4 28.1 1385.1
INTER (ours) 59.0 20.8 30.5 1595.5
LLaVA-v1.5 [38] M3ID [17] 67.1 19.7 30.9 1322.9
Ritual [59] 52.4 15.8 31.2 1754.7
SID  [27] 52.0 14.3 31.0 1692.4
INTER (ours) 51.8 14.1 31.9 1731.6
Table 3: Comparison with other methods.
@@ 5.3 Further Analysis
Comparison with Other Methods. We conducted experiments with M3ID [17], Ritual [59] and SID [27] in Tab. 3. The results demonstrate that our INTER achieves comparable performance among compared methods.
Performance on Other Types of Tasks or Different LVLMs. We conducted experiments with DeepSeek-VL2 [60] on the visual grounding task. As shown in Tab. 5, results show that the INTER boosts the model performance on this task.
Models RefCOCO [64]
testA testB
DeepSeek-VL2-Tiny [60] 87.8 78.4
+INTER 88.6 78.6
Table 5: Performance on other types of tasks or different LVLMs.
The Range of the Metric $\mathrm{I}(A)^{\textbf{y}_{t}}$. The value range of $\mathrm{I}(A)^{\textbf{y}_{t}}$ could be influenced by several factors, e.g., benchmarks, LVLMs, etc. These complexities make it challenging to establish a theoretical bound for its value range. Nevertheless, we conducted experiments to empirically assess the distribution of $\mathrm{I}(A)^{\textbf{y}_{t}}$ in Tab. 2. Moreover, when $\mathrm{I}(A)^{\textbf{y}_{t}}$ is negative, we consider that such interaction effects may hinder sampling this candidate token, which is considered similarly in prior studies [57, 70].

### v2

@@ Abstract
Hallucinations in large vision-language models (LVLMs) pose significant challenges for real-world applications, as LVLMs may generate responses that appear plausible yet remain inconsistent with the associated visual content. This issue rarely occurs in human cognition. We argue that this discrepancy arises from humans’ ability to effectively leverage multimodal interaction information in data samples. Specifically, humans typically first gather multimodal information, analyze the interactions across modalities for understanding, and then express their understanding through language. Motivated by this observation, we conduct extensive experiments on popular LVLMs and obtained insights that surprisingly reveal human-like, though less pronounced, cognitive behavior of LVLMs on multimodal samples. Building on these findings, we further propose INTER: Interaction Guidance Sampling, a novel training-free algorithm that mitigate hallucinations without requiring additional data. Specifically, INTER explicitly guides LVLMs to effectively reapply their understanding of multimodal interaction information when generating responses, thereby reducing potential hallucinations. On six benchmarks including VQA and image captioning tasks, INTER achieves an average improvement of up to 3.4% on five LVLMs compared to the state-of-the-art decoding strategy. The codes are released on Github.
@@ 3.5 Verifying the Impact of Interactions
It is challenging to quantify the relationship between the strength of multimodal interactions for keyword tokens and the overall accuracy of the response. To address this, we propose conducting experiments on the MME [19] benchmark for convenient verification. In MME, the keywords in the responses are limited to “yes" or “no", which are also used to determine the accuracy of the responses. This setup allows us to measure the strength of multimodal interactions for keyword tokens using $\mathrm{I}(A)_{r}^{\textbf{y}_{t}}$, where $\textbf{y}_{t}\in\{\textit{yes},\textit{no}\}$ and $A=\{\textbf{v},\textbf{p}\}$, for each response $r$. The appearance of the yes/no keywords directly indicates the binary correctness of the response. Furthermore, to better verify the trend of the positive impact of multimodal interactions, we propose to partition all the generated responses on MME into multiple bins based on the value of $\mathrm{I}(A)_{r}^{\textbf{y}_{t}}$ ($\textbf{y}_{t}\in\{yes,no\}$) sorted in ascending order, with each bin containing $Q$ samples. Then, we calculate the average contribution of interactions within each bin $\mathbf{b}$, denoted by $\overline{I}_{\mathbf{b}}=\frac{1}{Q}\sum_{r\in\mathbf{b}}\mathrm{I}(A)_{r}^{\textbf{y}_{t}}$. We denote $\overline{C}_{\mathbf{b}}$ as the ratio of correct responses in each bin $\mathbf{b}$. Thus, we can analyze the correlation between $\overline{I}_{\mathbf{b}}$ and $\overline{C}_{\mathbf{b}}$ among different bins, allowing us to verify trends in the positive impact of multimodal interactions on response accuracy.
@@ 5 Experiment
method InstructBLIP [13] LLaVA-v1.5 (7B) [38] Qwen-VL [3] mPLUG-owl2 [63]
COCO [35] AOKVQA [47] COCO [35] AOKVQA [47] COCO [35] AOKVQA [47] COCO [35] AOKVQA [47]
[…]
Table 2: The average F1-score on the POPE benchmark [34]. $\uparrow$ means that higher values indicate lower hallucination levels. The results indicate that applying INTER calibration, the models showed a reduction in hallucinations.
@@ 5.2 Experimental Results
model benchmark Nucleus Nucleus+INTER Beam Beam+INTER VCD* VCD*+INTER OPERA OPERA+INTER
LLaVA-v1.5 (7B)  [38] MM-Bench [39] $\uparrow$ 57.3 62.6 65.1 65.1 62.5 62.9 65.0 65.0
MMStar [7] $\uparrow$ 29.3 31.9 31.1 31.7 30.3 31.1 31.4 32.9
mPLUG-owl2  [63] MM-Bench [39] $\uparrow$ 57.0 61.4 63.5 63.7 59.2 59.5 63.4 63.6
MMStar [7] $\uparrow$ 30.5 32.3 30.7 30.5 31.5 32.3 30.5 30.9
Table 3: Validation of INTER on MM-Bench [39] and MMStar [7]. $\uparrow$ means that higher values indicate lower hallucination levels.
@@ 5.3 Further Analysis
Effectiveness at Different Parameter Scales. Our evaluation establishes INTER’s cross-scale adaptability through validation across scales spanning 1B to 26B parameters. As shown in Table 5, the MME [19] benchmark reveals consistent performance improvements across diverse scales, demonstrating scaling-agnostic generalization capabilities.
method InternVL2 [9, 10] LLaVA-v1.5 [38]
1B 4B 8B 26B 7B 13B
[…]
Table 5: Effectiveness of INTER on different parameter scales.
@@ 6 Conclusion
Acknowledgements. This work was supported by the National Natural Science Foundation of China under Grant No. U22A2098; the Key Science and Technology Development Plan of Jilin Province under Grant No. 20240302078GX. This work was supported by Alibaba Group through Alibaba Research Intern Program.